OpenStudy (anonymous):

6.02] and [6.03] Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

Please show me in elimination. :)

OpenStudy (anonymous):

For part 1 solve the first or second equation for any variable and plug it into the second equation.

OpenStudy (anonymous):

How? That is what I don't get. :(

OpenStudy (anonymous):

Would x be 6?

OpenStudy (anonymous):

You have the following equation: x - y = 16 2x + y = 2 Solve the second one for y y=-2x+2

OpenStudy (anonymous):

Plug this into the first equation

OpenStudy (anonymous):

-2x+2 x - (-2x+2) = 16

OpenStudy (anonymous):

x - (-2x+2) = 16 solve this for x.

OpenStudy (anonymous):

What is it? I don't understand. :(

OpenStudy (anonymous):

x=6, y=-10

OpenStudy (anonymous):

I was just showing you how to solve it using elimination.

OpenStudy (anonymous):

Oh, okay, thx, lol. :) :)

OpenStudy (anonymous):

Wait, the method I showed you was substitution, not elimination. My bad.

OpenStudy (anonymous):

Oh, well , either way is fine. :D